The following applies to a four-lane freeway. Determine LOS.
- Volume is 2,400 vph (one direction)
- Peak Hour Factor: 0.9
- BFFS = 60 mph
- 5% truck traffic
- 12-ft lanes
- 10-ft outside shoulders
- One interchange every 2 miles
- Level Terrain
- No recreational vehicles
- Commuter traffic
Answer #1

Base Free Flow Speed (BFFS) = 60 mph

Lane Width = 12 ft

Reduction of speed for Lane Width, f​​​​​​LW​​​​ = 0 mph

Lateral Clearance = 10 ft

Reduction of speed for Lateral Clearance, f​​​​​​LC = 0 mph

Interchanges = 1/ 2 mile = 0.5/mile

Reduction of speed based on Interchanges, f​​​​​IC = 0 mph

No. of lanes in each direction = 2

Reduction of speed for number of lanes, f​​​​​N = 4.5 mph

Free Flow Speed (FFS) = BFFS - f​​​​​​LW - f​​​​​​LC - f​​​​​N - fIC = 60 - 0 - 0 - 4.5 - 0 = 55.5 mph

Peak Flow, V = 2400 veh/hr

Peak-hour factor = 0.9

Trucks = 5 %

Level Terrain

fHV = 1/ (1 + 0.05 (1.5-1)) = 1/1.025 = 0.976

fP = 1.0

Peak Flow Rate, Vp = V / (PHV*n*fHV*fP) = 2400/ (0.90*2*0.976*1.0) = 1366.12 ~ 1367 veh/hr/ln

S = FFS

S = 55.5 mph

Density = Vp/S = (1367) / (55.5) = 24.63 veh/mi/ln

LOS C

Density of LOS C should lie between 18 – 26 veh/mi/ln